OpenStudy (anonymous):

prove 1+sin2x=(sinx=cosx)^2

OpenStudy (zarkon):

got a lot of equal signs there ;)

OpenStudy (zarkon):

for \[1+\sin(2x)=(\sin(x)+\cos(x))^2\] use \[\sin(2x)=2\sin(x)\cos(x)\] and \[\sin^2(x)+\cos^2(x)=1\]

OpenStudy (anonymous):

\[(\sin x + \cos x)^2 = \sin^2 x + 2(\sin x \cos x) + \cos^2 x\]

OpenStudy (anonymous):

zarkon: haha thanks for noticing that ;) and....? is that it?

OpenStudy (anonymous):

Nah you still have to work it out, just fit the pieces together lol

OpenStudy (zarkon):

combine the work Will! and I did.

OpenStudy (anonymous):

i don't get it :(

OpenStudy (zarkon):

\[(\sin x + \cos x)^2 = \sin^2 x + 2(\sin x \cos x) + \cos^2 x\] \[= \sin^2 x + \cos^2 x +2\sin x \cos x=\cdots\]

OpenStudy (zarkon):

now use what I told you to use. :)

OpenStudy (anonymous):

aha! solved! :) gracias :)

OpenStudy (zarkon):

De nada